OpenStudy (anonymous):

Medal and Fan Evaluate y = x – 13 when x = –3.

OpenStudy (anonymous):

@Michele_Laino help please

OpenStudy (secret-ninja):

When x = -3, all you have to do is substitute it for x and solve. Like this: \[y=-3-13\]

OpenStudy (igreen):

Just plug in -3 for 'x': y = x - 13 y = -3 - 13 Can you solve that?

OpenStudy (secret-ninja):

Solve it and you have your answer.

OpenStudy (igreen):

My comments are lagging.

OpenStudy (anonymous):

thank you i fan both medal 1

OpenStudy (michele_laino):

please you have to set x=-3 into your formula, namely: \[y=-3-13=...\] please complete the computation

OpenStudy (secret-ninja):

Mine too lol. Comments keep getting glitched these days.

OpenStudy (igreen):

Yep ^^

OpenStudy (secret-ninja):

No problem. :) @boygame20

OpenStudy (igreen):

Yay no more laggy comments :D